# REF-VLM: Triplet-Based Referring Paradigm for Unified Visual Decoding

### v1

@@ Abstract
Multimodal Large Language Models (MLLMs) demonstrate robust zero-shot capabilities across diverse vision-language tasks after training on mega-scale datasets. However, dense prediction tasks, such as semantic segmentation and keypoint detection, pose significant challenges for MLLMs when represented solely as text outputs. Simultaneously, current MLLMs utilizing latent embeddings for visual task decoding generally demonstrate limited adaptability to both multi-task learning and multi-granularity scenarios. In this work, we present REF-VLM, an end-to-end framework for unified training of various visual decoding tasks. To address complex visual decoding scenarios, we introduce the Triplet-Based Referring Paradigm (TRP), which explicitly decouples three critical dimensions in visual decoding tasks through a triplet structure: concepts, decoding types, and targets. TRP employs symbolic delimiters to enforce structured representation learning, enhancing the parsability and interpretability of model outputs. Additionally, we construct Visual-Task Instruction Following Dataset (VT-Instruct), a large-scale multi-task dataset containing over 100 million multimodal dialogue samples across 25 task types. Beyond text inputs and outputs, VT-Instruct incorporates various visual prompts such as point, box, scribble, and mask, and generates outputs composed of text and visual units like box, keypoint, depth and mask. The combination of different visual prompts and visual units generates a wide variety of task types, expanding the applicability of REF-VLM significantly. Both qualitative and quantitative experiments demonstrate that our REF-VLM outperforms other MLLMs across a variety of standard benchmarks. The code, dataset, and demo available at https://github.com/MacavityT/REF-VLM.
Figure 1: Overview of Visual Tasks Supported by REF-VLM. REF-VLM supports a wide range of visual tasks with user-provided visual inputs such as points, boxes, scribbles, and masks, while enabling the decoding of visual contents into formats like points, boxes and masks. For better visualization, details such as TRP and part of special tokens are hidden in the model’s responses.
Table 1: Comparisons of recent MLLMs and their capabilities in performing downstream tasks.
Visual
Understanding
Referring
Expression
Interactive
Grounding (IG)
Grounded Conversation
Generation (GCG)
Open Vocabulary
Identification
Model End-to-End Extensible VQA Caption RES REC REG Mask Box Mask Box OVS OVD FOVS FOVD
LLaVA [34] - - - - - - - - - - - -
MM-REACT [64] - - - - - - - - - -
Visual ChatGPT [54] - - - - -
HuggingGPT [48] - - - - -
BuboGPT [74] - - - - - - - - - - -
Kosmos-2 [41] - - - - - - - - -
Shikra [9] - - - - - - - - -
MiniGPT-v2 [8] - - - - - - - - -
NExT-Chat [70] - - - - - - -
Ferret [66] - - - - - - - -
SHPINX [31] - - - - - - - - -
LLaVA-Plus [35] - - - - - - - - -
LISA [28] - - - - - - - - - - -
Osprey [68] - - - - - - - - - - -
GLaMM [45] - - - - - - - - -
PixelLM [61] - - - - - - - - - -
PSALM [73] - - - - - - -
GroundHOG [72] - - - - - - - -
F-LLM [58] - - - - - - - - - -
VITRON [19] - - - - - - - - - -
VisionLLM [52] - - - - -
VisionLLMv2 [56] - - -
REF-VLM (Ours)
@@
1 Introduction
Multimodal Large Language Models (MLLMs) demonstrate excellent performance in tasks such as visual question answering and scene understanding [34, 2, 50]. Despite these achievements, typical MLLMs primarily understand input and generate responses with text, which limits their ability to perform fine-grained visual localization. As a result, they struggle to make significant contributions in real-world applications such as autonomous driving, robotics, and medical diagnosis.
In this work, we present REF-VLM, an end-to-end framework that enables unified multi-task training, contrasting with existing models that require separate stage-wise training for different tasks, thereby enhancing semantic consistency. To encode diverse user interactions, we introduce a novel parameter-free Mask-Guided Aggregation scheme. Additionally, we propose a Latent Embeddings Router and Parallel Group Hungarian Matching to handle multi-task and multi-granularity decoding scenarios effectively.
As illustrated in Figure 2 (b), conventional Vision Language Models (VLMs) [45, 28, 56] typically generate responses in a simplistic “Visual Concept + Referring token” format, which proves inadequate for complex multi-granularity scenarios. To address even more challenging visual decoding tasks as shown in Figure 2 (c), we introduce the Triplet-Based Referring Paradigm (TRP), which enforces explicit generation of three core components: (1) visual concept, (2) decoding type, and (3) referring tokens, organized through a structured special token framework. The inherent compositional nature of the triplet structure endows TRP with a “one-fits-all” capability, effectively handling complex multi-task and multi-granularity scenarios. To further enhance the precision of structured text generation, we introduce Visual Decoding Chain-of-Thought (VD-CoT), which requires the model to first overview the image and summarize task-relevant information before generating TPR-compliant responses. The synergistic combination of VD-CoT and TRP significantly improves model performance and accuracy across multi-task visual decoding scenarios.
To enhance the diversity of vision-language tasks, we propose Visual-Task Instruction Following Dataset (VT-Instruct), a multimodal dataset specifically designed to support a wide range of tasks, including Visual Understanding [34], Referring Expressions [9, 66, 70, 73], Interactive Grounding (IG) [73, 56], Open-Vocabulary Identification [56, 73, 54, 48], Grounded Conversation Generation (GCG) [45], Keypoint Detection [56, 31] and Depth Estimation [31]. VT-Instruct consists of more than 100 million high-quality multimodal dialogue samples, primarily derived from publicly available datasets such as LAION-5B [47], SA-1B [26], COCO [30], GRIT [41], etc. Each sample is enhanced with thoughtfully crafted prompt templates with multimodal inputs (e.g. images, texts, points, boxes, scribbles and masks) to facilitate instruction following and diverse outputs (e.g. texts, boxes, keypoints, depth and masks) for different downstream tasks.
Our contributions can be summarized as follows:
  • We propose REF-VLM, an end-to-end framework for unified visual decoding tasks, integrating novel components like Mask-Guided Aggregation, Latent Embeddings Router, and Parallel Group Hungarian Matching to boost multi-task performance and adaptability.
  • We design a Unified Instruction Pipeline with the Triplet-Based Referring Paradigm and VD-CoT for precise referring in multi-task and multi-granularity scenarios. We also introduce VT-Instruct, a large-scale dataset with 100M multimodal dialogue samples spanning 25 task types, enabling robust understanding and decoding of diverse visual units.
  • Extensive experiments show that REF-VLM outperforms existing MLLMs on tasks including Visual Understanding, Referring Expression, Grounded Conversational Generation, Open-Vocabulary Identification, and Interactive Grounding.
2 Related Works

### v2

@@ Abstract
Multimodal Large Language Models (MLLMs) demonstrate robust zero-shot capabilities across diverse vision-language tasks after training on mega-scale datasets. However, dense prediction tasks, such as semantic segmentation and keypoint detection, pose significant challenges for MLLMs when represented solely as text outputs. Simultaneously, current MLLMs utilizing latent embeddings for visual task decoding generally demonstrate limited adaptability to both multi-task learning and multi-granularity scenarios. In this work, we present REF-VLM, an end-to-end framework for unified training of various visual decoding tasks. To address complex visual decoding scenarios, we introduce the Triplet-Based Referring Paradigm (TRP), which explicitly decouples three critical dimensions in visual decoding tasks through a triplet structure: concepts, decoding types, and targets. TRP employs symbolic delimiters to enforce structured representation learning, enhancing the parsability and interpretability of model outputs. Additionally, we construct Visual-Task Instruction Following Dataset (VT-Instruct), a large-scale multi-task dataset containing over 100 million multimodal dialogue samples across 25 task types. Beyond text inputs and outputs, VT-Instruct incorporates various visual prompts such as point, box, scribble, and mask, and generates outputs composed of text and visual units like box, keypoint, depth and mask. The combination of different visual prompts and visual units generates a wide variety of task types, expanding the applicability of REF-VLM significantly. Both qualitative and quantitative experiments demonstrate that our REF-VLM outperforms other MLLMs across a variety of standard benchmarks. The code, dataset, and demo will be publicly available.
I Introduction
Multimodal Large Language Models (MLLMs) show strong performance in visual question answering and scene understanding [33, 2, 49, 51], yet their text-centric design limits fine-grained visual units decoding, hindering practical deployment in domains such as autonomous driving, robotics, and medical diagnosis.
As illustrated in Fig. 1(b), existing vision–language models (VLMs) [43, 28, 54] typically decode visual units (e.g., bounding boxes and masks) using a simplistic “Visual Concept + Referring Token” format, which is insufficient for complex, multi-granularity scenarios. To tackle more challenging visual decoding tasks (Fig. 1(c)), we propose the Triplet-Based Referring Paradigm (TRP), which explicitly generates three components—(1) visual concept, (2) decoding type, and (3) referring tokens—via a structured special-token framework. Owing to its compositional triplet design, TRP provides a “one-fits-all” capability for unified multi-task and multi-granularity visual decoding.
To enhance the diversity of vision–language tasks, we introduce the Visual-Task Instruction Following Dataset (VT-Instruct), a large-scale multimodal dataset designed to support a broad range of tasks, including visual understanding [33], referring expressions [23, 8, 62, 66, 69], interactive grounding (IG) [69, 54], open-vocabulary identification [59, 54, 69, 52, 46], grounded conversational generation (GCG) [43], keypoint detection [54, 31], and depth estimation [31]. VT-Instruct contains over 100M high-quality multimodal dialogue samples curated primarily from publicly available datasets such as LAION-5B [45], SA-1B [26], COCO [30], and GRIT [39]. Each sample is augmented with carefully designed prompt templates involving multimodal inputs (e.g., images, text, points, boxes, scribbles, and masks) to support instruction following and diverse task-specific outputs.
Building upon the proposed decoding paradigm and the constructed dataset, we present REF-VLM, an end-to-end framework that enables unified multi-task training, contrasting with existing models that require separate stage-wise training for different tasks, thereby enhancing semantic consistency. To encode diverse user interactions, we introduce a novel parameter-free Mask-Guided Aggregation scheme. Additionally, we propose a Latent Embeddings Router and Parallel Grouped Hungarian Matching to to enable the integration of multiple visual unit decoders in a joint multi-task training framework. Moreover, its plug-and-play architecture provides enhanced extensibility.
Our contributions can be summarized as follows:
  • We introduce a Triplet-Based Referring Paradigm (TRP) for precise referring under multi-task and multi-granularity settings. To support large-scale training under this paradigm, we construct VT-Instruct, a large-scale multimodal dataset comprising 100M dialogue samples across 25 visual task types, enabling robust understanding and decoding of diverse visual units.
  • We propose REF-VLM, an end-to-end framework for unified visual decoding, integrating novel components such as Mask-Guided Aggregation, Latent Embeddings Router, and Parallel Grouped Hungarian Matching to improve multi-task performance and adaptability.
  • Extensive experiments show that REF-VLM consistently outperforms existing MLLMs across diverse tasks, including visual understanding, referring expression comprehension, grounded conversational generation, open-vocabulary identification, and interactive grounding.
@@ I Introduction
II Related Works
@@
Decode Visual Units as Sequences. The most straightforward solution leverages the text generation capabilities of MLLM to produce visual localization results in textual format [8, 9, 39, 37]. This approach does not require structural modifications to the MLLM. Pix2Seq [9] leverages the model’s autoregressive generation capability to express bounding boxes and class labels as sequences of discrete tokens. Shikra [8] constructs an visual supervised fine-tuning dataset, where the model needs to perform inductive analysis before answering complex questions, and outputs bounding boxes in text form.